In a stack there are 5 books each of thickness 20 mm and 5 paper sheets each of thickness 0.016 mm. What is the total thickness of the stack?

Grade:12th pass

2 Answers

Pawan Prajapati
2 years ago
In this question first calculate the thickness of 5 books and then thickness of 5 sheets with the given data. Hence, the total thickness of the stack is the sum of thickness of 5 books and the thickness of 5 sheets. So, use this concept to reach the solution of the problem. Complete step-by-step answer: Given, Number of books = 5 Thickness of each book = 20 mm So, thickness of 5 books =5×20=100 mm Number of paper sheets = 5 Thickness of each paper sheet = 0.016 mm So, thickness of 5 paper sheets =5×0.016=0.080 mm Here total thickness of the stack = thickness of 5 books + thickness 5 papers = 100 mm + 0.080 mm = 100.08 mm Therefore, the total thickness of the stack is 100.08 mm or 1.0008 cm Note: We can write the total thickness of the stack in terms of millimeters (mm) or in terms of centimeters (cm). To convert millimeters (mm) into centimeters (cm) we use the conversion 100 mm=1 cm .
Ayush 8A 7239
2 years ago
Thickness of 1 book = 20mm
Thickness of 5 books = 5 × 20 mm = 100 mm
 
Thickness of 1 paper sheet = 0.016 mm
Thickness of 5 paper sheets = 5 × 0.016 = 0.08 mm
 
Total Thickness of the stack = 100 mm + 0.08 mm = 100.08 mm or 1.0008 × 102
 
Hope it helped you